OpenStudy (anonymous):

Solve the equation and check your solutions. If there is no solution, say so. 10/3x-8 = 2/x

OpenStudy (anonymous):

\[\frac{10}{3x-8}=\frac{2}{x}.......2(3x-8)=10x........6x-16=10x.......4x=-16....x=-4\]